Multiply out (3x3 + 4x - 5)(2x3 - 8x2 + 2) using polynomial long multiplication.


1
Expert's answer
2022-04-11T15:02:38-0400

"(3x^3+4x-5)(2x^3-8x^2+2)=\\=
3x^3\cdot2x^3+3x^3\cdot(-8x^2)+3x^3\cdot2+\\
+4x\cdot2x^3+4x\cdot(-8x^2)+4x\cdot2-\\
-5\cdot2x^3-5\cdot(-8x^2)-5\cdot2=\\
=6x^6-24x^5+6x^3+\\
+8x^4-32x^3+8x-\\
-10x^3+40x^2-10=\\
=6x^6-24x^5+8x^4-36x^3+40x^2+8x-10"